Circular No. 6407

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SATELLITES OF SATURN F. Roddier, Institute for Astronomy, University of Hawaii (IfA); A.

Brahic, Observatoire de Paris and Université de Paris (Paris); C. Dumas, IfA; C. Ferrari, Paris; J. E. Graves, M. J. Northcott, and T. Owen, IfA; L. Perret, Paris; C. Roddier, IfA; and P. Thebault, Paris, report: “A new adaptive-optics system developed at the University of Hawaii was operated at the 3.6-m Canada-France-Hawaii telescope during four nights spanning the 1995 Aug. ring-plane crossing. Several hundred 15-, 30-, and 60-s exposures were obtained in the bands I, J, H, and K; the average angular resolution is 0′′.15 in the H band. Saturn X (Janus), XI (Epimetheus), XII (Helène), XIII (Telesto), XIV (Calypso), XV (Pandora), and XVI (Prometheus) were identified; positions are in fair agreement with the new ephemerides deduced from Hubble Space Telescope (HST) observations. Three additional objects — S/1995 S 8, S/1995 S 9, and S/1995 S 10 — have been discovered. The number of fitted observations is given below, followed by our estimate of the equivalent orbital radius and longitude for epoch 1995 Aug. 10.5 TDT (at Saturn), measured from the ascending node of Saturn’s equatorial plane on the earth’s J2000.0 equator; the fits include the effects of Saturnian J_{2} and J_{4} harmonics: S/1995 S 8, 16 observations, 141 400 ± 1000 km, 323^{o} ± 1^{o}; S/1995 S 9, 16, 141 400 ± 2000 km, 315^{o} ± 1^{o}.8; S/1995 S 10, 8, 140 050 ± 100 km, 131^{o}.4 ± 1^{o}.1. S/1995 S 10 has been observed on both Aug. 9 and 10; it is co-orbital to the F ring. This object is most probably the candidate satellite S/1995 S 5, which was observed by the HST on Aug. 10. S/1995 S 8 may be identified with S/1995 S 7 observed by HST, since the observations are consistent. The S/1995 S 9 azimuthal extent of 5^{o} suggests an arc structure, which may be embedded in the F ring; it cannot be linked with any other objects discovered by HST at the same epoch.”

1996 JA_{1}

Further to IAUC 6402, M. Hicks reports that U. Fink was also involved with the observations mentioned there. He adds that the second-night data, which were obtained with the spectrograph slit widened to 10′′ in order to eliminate tracking errors and differential refraction, are in complete agreement with the first-night data. The identification of 1996 JA_{1} as V type (eucrite) is thus very secure.

1996 May 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green